OpenStudy (anonymous):

Help Please! The equation of the line with the slope 1/2 and passing through (3,7) is?

OpenStudy (anonymous):

(y - y1) = m(x - x1) x1,y1 is the point they gave you m = slope they gave you that also plug in and re-write into y = mx+b

OpenStudy (anonymous):

Alright Thanks!